# Simulating stochastic population dynamics: The Linear Noise Approximation can capture non-linear phenomena

### v3

@@ 6 Discussion
The results in this paper demonstrate the capability of the modified LNA method to effectively capture non-linear dynamics over long time intervals. The accuracy of the approximation combined with the analytical tractability of the LNA method, and speed of computation establishes a robust framework for investigating population dynamics, and fitting these models to data. It overcomes key limitations of existing methods and opens new avenues for advancing the stochastic modeling of dynamical systems.
We showcase the versatility of our methodology across a range of systems exhibiting two fundamental forms of non-linear behavior—oscillations and bistability—which are common in biology and other scientific domains. Beyond these cases, our approach, grounded in center manifold theory from dynamical systems, is applicable to a much broader class of reaction networks. Specifically, for any network whose RRE solution features a non-hyperbolic equilibrium—a scenario frequently encountered in practice—the pcLNA algorithm enables fast and accurate simulation of long-term stochastic trajectories, even in large and complex systems.
@@
To simplify notation, in this section we use the notation $x_{t}=x(t)$, for all states considered (i.e. $\bm{x}$}, $\bm{\xi}$, $\bm{X}$}). The LNA ansantz,
${\bm{X}}(t)={\bm{x}}(t)+\Omega^{-1/2}\bm{\xi}(t),$
@@
$\sqrt{\Omega}\left(\rho_{j}({\bm{X}}_{t})-\rho_{j}({\bm{x}}_{t})\right)=\bm{\xi}(t)^{T}\nabla_{\!{\bm{x}}}\,\rho_{j}({\bm{x}}_{t})+O\left(\Omega^{-1/2}\right),$ (23)
where $\nabla_{\!{\bm{x}}}^{T}=(\partial/\partial x_{1},\dots,\partial/\partial x_{n})$. Hence, for sufficiently large $\Omega$,
$\bm{\xi}_{t+dt}-\bm{\xi}_{t}\approx\sum_{j}\nu_{j}\bm{\xi}(t)^{T}\nabla_{\!{\bm{x}}}\,\rho_{j}({\bm{x}}_{t})dt+\sum_{j}Z_{j}\nu_{j}\sqrt{\rho_{j}({\bm{x}}_{t})dt}$
@@
$d\bm{\xi}=\bm{J}\bm{\xi}dt+{\bm{E}}\,d\!{\bm{B}}_{t},$ (24)
where $\bm{J}$ is the Jacobian matrix of the RRE, ${\bm{E}}={\bm{E}}({\bm{x}})={\bm{S}}\text{diag}\left(\!\sqrt{\rho_{1}({\bm{x}})},\dots,\sqrt{\rho_{r}({\bm{x}})}\right)$ the product of the stoichiometry matrix ${\bm{S}}=[{\bm{\nu}}_{1}\cdots{\bm{\nu}}_{r}]$, and the square root of the diagonal matrix, with main diagonal entries $(\rho_{1}({\bm{x}}),\dots,\rho_{r}({\bm{x}}))$, and ${\bm{B}}_{t}$ an $n$-dimensional Wiener process.
@@
Consider two ODE systems as in 6 with velocities, $\bm{F}$} and $\bm{F}$$^{\prime}$, on $\mathbb{R}^{n}$ and $\mathbb{R}^{m}$ respectively. Suppose they each have an equilibrium point, ${\bm{x}}_{0}$ and ${\bm{x}}_{0}^{\prime}$, respectively. Then, near their equilibrium, the two dynamical systems are locally topologically equivalent if there exists a homeomorphism ${\bm{h}}:\mathbb{R}^{n}\to\mathbb{R}^{m}$ that is,
@@
    maps paths of the system with velocity $\bm{F}$} in the neighbourhood $A$ onto paths of the system with velocity $\bm{F}$$^{\prime}$ in the neighbourhood ${\bm{h}}(A)\subset\mathbb{R}^{m}$, preserving the direction of time.
Intuitively, this means the paths of one dynamical system near its equilibrium can be continuously deformed onto paths of the other dynamical system near its equilibrium, and hence both systems exhibit paths that are qualitatively similar, i.e. have the same key features, near their equilibria.
The deterministic evolution $\{{\bm{x}}(t)\ |\ t\geq 0\}$ solving the reaction rate equation (RRE) (6) when one parameter, say $\alpha\in\mathbb{R}$, is free to vary. We assume that the system has a non-hyperbolic equilibrium. That is, there exists an $\alpha_{p}\in\mathbb{R}$ such that,
@@
where $n_{0},n_{+},n_{-}$ denote the number of eigenvalues of the Jacobian matrix $\bm{J}({\bm{x}}_{\alpha_{p}},\alpha_{p})$ with zero, positive and negative real part, respectively. Following this, consider the extended system,
$\left\{\begin{aligned} \dot{{\bm{x}}}&={\bm{F}}({\bm{x}},\alpha),\\ \dot{\alpha}&=0.\end{aligned}\right$. (26)
Note that the Jacobian matrix of (26) evaluated at ${\bm{x}}={\bm{x}}_{\alpha_{p}}$ and $\alpha=\alpha_{p}$ is the $(n+1)\times(n+1)$ matrix,
$\begin{pmatrix}0&0\\ \left.\frac{\partial{\bm{F}}}{\partial\alpha}\right|_{\begin{subarray}{c}{\bm{x}}={\bm{x}}_{\alpha_{p}}\\ \alpha=\alpha_{p}\end{subarray}}&\bm{J}({\bm{x}}_{\alpha_{p}},\alpha_{p})\end{pmatrix}$, (27)
with $n_{0}+1$ eigenvalues with zero real part and $n-n_{0}$ eigenvalues with nonzero real part. Under the assumptions above, the following theorem holds (see [34, Theorem 5.1 and Lemma 5.1]).

### v4

@@ 7 Discussion
The results in this paper demonstrate the capability of the modified LNA method to effectively capture non-linear dynamics over long time intervals. The modification introduced is justified by a novel reformulation of the LNA in coordinates that separate persistent non-linear dynamics from transient dynamics. The accuracy of the approximation combined with the analytical tractability of the LNA method, and speed of computation establishes a robust framework for investigating population dynamics, and fitting these models to data. It overcomes key limitations of existing methods and opens new avenues for advancing the stochastic modeling of dynamical systems.
We showcase the the versatility of our methodology across a range of systems exhibiting two fundamental forms of non-linear behavior—oscillations and bistability—which are common in biology and other scientific domains. Beyond these cases, our approach, grounded in centre manifold theory from dynamical systems, is applicable to a much broader class of reaction networks. Specifically, for any parameter-dependent network whose RRE solution features a non-hyperbolic equilibrium—a scenario frequently encountered in practice—the pcLNA algorithm enables fast and accurate simulation of long-term stochastic trajectories, even in large and complex systems.
@@
The following is a heuristic derivation of the LNA equations, similar to this provided in [1]. For a formal derivation see [36]. To simplify notation, in this section we use the notation $x_{t}=x(t)$, for all states considered (i.e. $\bm{x}$}, $\bm{\xi}$, $\bm{X}$}). The LNA ansantz,
${\bm{X}}_{t}={\bm{x}}_{t}+\Omega^{-1/2}\bm{\xi}_{t},$
@@
$\sqrt{\Omega}\left(\rho_{j}({\bm{X}}_{t})-\rho_{j}({\bm{x}}_{t})\right)=\bm{\xi}(t)^{\top}\nabla_{\!{\bm{x}}}\,\rho_{j}({\bm{x}}_{t})+O\left(\Omega^{-1/2}\right),$ (37)
where $\nabla_{\!{\bm{x}}}^{\top}=(\partial/\partial x_{1},\dots,\partial/\partial x_{n})$. Hence, for sufficiently large $\Omega$,
$\bm{\xi}_{t+dt}-\bm{\xi}_{t}\approx\sum_{j}\lambda_{j}\bm{\xi}(t)^{\top}\nabla_{\!{\bm{x}}}\,\rho_{j}({\bm{x}}_{t})dt+\sum_{j}Z_{j}\lambda_{j}\sqrt{\rho_{j}({\bm{x}}_{t})dt}$
@@
$d\bm{\xi}_{t}=\bm{J}_{t}\bm{\xi}_{t}dt+{\bm{E}}_{t}\,d\!{\bm{B}}_{t},$ (38)
where $\bm{J}_{t}=\bm{J}({\bm{x}}_{t})$ is the Jacobian matrix of the RRE (6), and
${\bm{E}}_{t}={\bm{E}}({\bm{x}}_{t})={\bm{S}}\text{diag}\left(\!\sqrt{\rho_{1}({\bm{x}}_{t})},\dots,\sqrt{\rho_{r}({\bm{x}}_{t})}\right)$
the product of the stoichiometry matrix ${\bm{S}}=[{\bm{\nu}}_{1}\cdots{\bm{\nu}}_{r}]$, and the square root of the diagonal matrix, with main diagonal entries $(\rho_{1}({\bm{x}}),\dots,\rho_{r}({\bm{x}}))$, while ${\bm{B}}_{t}$ an $n$-dimensional Wiener process.
@@
For a more detailed introduction to centre manifold theory see [37].
@@
Consider two ODE systems as in 6 with velocities $\bm{F}$} and $\bm{F}$$^{\prime}$ on $\mathbb{R}^{n}$ and $\mathbb{R}^{m}$ respectively. Suppose they each have an equilibrium point ${\bm{x}}_{0}$ and ${\bm{x}}_{0}^{\prime}$, respectively. Then, near their equilibrium, the two dynamical systems are locally topologically equivalent if there exists a homeomorphism ${\bm{h}}:\mathbb{R}^{n}\to\mathbb{R}^{m}$ that is
@@
    maps paths of the system with velocity $\bm{F}$} in $A$ onto paths of the system with velocity $\bm{F}$$^{\prime}$ in ${\bm{h}}(A)\subset\mathbb{R}^{m}$, preserving the direction of time.
Intuitively, this means the paths of one dynamical system near its equilibrium can be continuously deformed onto paths of the other dynamical system near its equilibrium, and hence both systems exhibit qualitatively similar behavior near their equilibria.
Extended system and non-hyperbolic equilibrium
Consider the deterministic evolution solving the reaction rate equation (6) when one parameter $\alpha\in\mathbb{R}$ is free to vary. Assume there exists $\alpha_{p}\in\mathbb{R}$ such that
@@
where $n_{0},n_{+},n_{-}$ denote the number of eigenvalues of the Jacobian matrix $\bm{J}({\bm{x}}_{\alpha_{p}},\alpha_{p})$ with zero, positive and negative real part, respectively.
Introduce the extended autonomous system
$\left\{\begin{aligned} \dot{{\bm{x}}}&={\bm{F}}({\bm{x}},\alpha),\\ \dot{\alpha}&=0,\end{aligned}\right$. (40)
whose Jacobian at $({\bm{x}},\alpha)=({\bm{x}}_{\alpha_{p}},\alpha_{p})$ is
$\mathcal{A}:=\begin{pmatrix}\bm{J}({\bm{x}}_{\alpha_{p}},\alpha_{p})&\partial_{\alpha}{\bm{F}}({\bm{x}}_{\alpha_{p}},\alpha_{p})\\ 0&0\end{pmatrix}$. (41)
This matrix has $n_{0}+1$ eigenvalues with zero real part and $n-n_{0}$ eigenvalues with nonzero real part.